Reports of 17321
VIS (General)
takafumi.ushiba - 19:43 Wednesday 08 December 2021 (19133) Print this report
Automesurement: ITMX and ETMX

Automesurement of suspension plant is running for ITMX and ETMX tonight.
So, please do not touch these suspension until tomorrow morning.

REFNUMs are 0010 and 0007 for ITMX and ETMX, respectively.

CRY (Cryo-payload EX)
takafumi.ushiba - 16:36 Wednesday 08 December 2021 (19132) Print this report
Comment to EX payload finalization Day4: final visual inspection of ETMX (19061)

[Tamaki, Ushiba]

We went to Xend and checked moving mass position.
There are enough space between moving mass and their stoppers.
Photos are stored at following links:
https://www.dropbox.com/sh/r5e4d6vlgan327n/AAAdFb8BOvGzjsuQzXRzHRpfa?dl=0

Also, we checked the OpLev beam path after setting SIs around radiation shields.
However, a part of SIs are very close to beam path and it is hard to recover them today.
So, I would like to go to Xend again with Kimura-san and optimize them tomorrow or the day after tomorrow.
Photos are stored at following links:
https://www.dropbox.com/sh/2mia1jtoegf5wf2/AACqK6Tiq48PGT7O0Tz2W7t3a?dl=0

VIS (PR3)
hiromi.yasui - 16:17 Wednesday 08 December 2021 (19131) Print this report
Inspection of mirror

With Hirata-san,

We started taking pictures to determine whether mirrors should be cleaned or not.

Today, we tried taking some pictures of the PR3 HR side.

We will continue to take better pictures under different conditions with different light positions and brightness.

AOS (Optical Levers)
takaaki.yokozawa - 14:11 Wednesday 08 December 2021 (19129) Print this report
Trial of oplev beam alignment for PRM and PR2
[K.Tanaka, Yuzu, Yokozawa]

We performed the oplev beam alignment for PRM and PR2.

For PRM, we adjust to hit the oplev beam to PRM centering.(Fig.1.)
But we didn't know the angle of PRM, we didn't touch the receiver optics.
During this work, we unfortunately touched the suspension by IR card...
We just set the PSD on the optical table of PRM oplev receiver port.

For PR2, because we didn't touch anything because we couldn't judge that we can open the chamber of +X side.
(And heavy rain.)
Images attached to this report
VAC (BS)
shinji.miyoki - 13:33 Wednesday 08 December 2021 (19130) Print this report
No lost item in EXA tank

[miyoki, chendan]

We inspect the inside of the EXA tank (actually the area above the optical table) to make assure there is no nasty thing inside EXA. According to our eyeball inspection. no nasty thing inside. 

(After O3GK, several tapes were found in this tank.)

VIS (IMM)
takaaki.yokozawa - 12:50 Wednesday 08 December 2021 (19128) Print this report
Comment to Check the optics for the IMMT1 oplev (19115)
One day trend (Sum value was about 1,100 from Yesterday, but we didn't perform centering yesterday...)
Images attached to this comment
MIF (General)
kenta.tanaka - 12:49 Wednesday 08 December 2021 (19127) Print this report
We found the beam profiler for large beam size (~70 mm)

Yokozawa, Yuzurihara, K. Tanaka

We found the beam profiler for large beam size (~70mm) in the JGW doc (https://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=872). (fig, 1)

 we tried to connect it to my PC and check the operation, but the RayCi software did not recognize the profiler, so we gave up today. Yokozawa-san will investigate in detail later.

Now, we put the profiler in the pre-room of PR clean booth. (fig.2)

Images attached to this report
VIS (BS)
fabian.arellano - 12:18 Wednesday 08 December 2021 (19124) Print this report
Comment to IP ballast mass adjustment and resonant frequencies measurement. (16858)

Summary: the resonant frequencies in IP-L and IP-T have changed since the last time we adjueted them with ballast masses, however, there is no obvious reason.

Recently Terrence reported that the IP was very soft (entry 19089). He measured the resonant frequency in L to be 47 mHz. However, the latest time we adjusted it using ballast masse, it was 63 mHz (entry 17191).  He did his assessment with a temperature of 32.7 °C, almost at the same temperature of 23.8 °C we had when we adjusted such a frequency. Therefore, unless the thermometer (ondotori) has been malfunctioning , this time I cannot blame the temperature for this change.

I measured transfer functions again today (see plots). The following observations hold:

  • Resonant frequencies in L and T respectively: 50, 82 mHz.
  • Values reported in entry 17191: 63, 92 mHz.
  • Frequency resolution: 5 mHz.
  • Temperature: 23.5 °C.
  • I moved the IP close to the setpoint before measuring the TFs.

There are similarities and differences:

  • In both transfer functions there is low coherence between 160 and 380 mHz whereras it is higher in the TFs reported in entry 17191.
  • In the IP-L TF, at low frequencies, the motion in T is 180° out of phase with respect to L, whereas it is in phase in the TF reported in entry  17191.
  • The difference between the resonance frequencies in L and T is 32 mHz, similar to the 29 mHz reported in entry  17191.

Currently, I don't know what could have produced the change.

Images attached to this comment
VIS (EX)
satoru.ikeda - 11:22 Wednesday 08 December 2021 (19126) Print this report
Comment to OFFLOAD of EX (19098)

I checked the past movement amount and the movement direction of KeyStone.
I also used this information to adjust the orientation of the buttons in OFFLOAD_OVERVIEW.adl.

F0: + step > keyston position up, -51200step = -123.8um at VIS-ETMX_F0_LVDTINF_GAS_OUT16
 => about +30[um]/12800[cnt]
F1: + step > keyston position down, -32000step = +256.0um at VIS-ETMX_F1_LVDTINF_GAS_OUT16
 => about -100[um]/12800[cnt]
F2: Not used.(K-Log#17715)
F3: + step > keyston position down, +6400step = -78.0um at VIS-ETMX_F1_LVDTINF_GAS_OUT16
 => about -160[um]/12800[cnt]
BF: + step > keyston position down, -32000step = -251.0um at VIS-ETMX_BF_LVDTINF_GAS_OUT16
 => about -100[um]/12800[cnt].

Images attached to this comment
Non-image files attached to this comment
VIS (IX)
ryutaro.takahashi - 11:19 Wednesday 08 December 2021 (19125) Print this report
Comment to ITMX GAS LVDT signals seemed strange. (18789)

[Takahashi, Sato]

We replaced the main card in the LVDT driver for the GAS filters (F1, F2, F3, BF) and the IP (F0, IP). The spectra were improved! The noise level was reduced very much. The F0 spectrum was also improved. This means that both outputs for the GASF and the IP were noisy in the original card.

Images attached to this comment
VIS (IX)
masahide.tamaki - 9:03 Wednesday 08 December 2021 (19123) Print this report
Comment to Health check (checking the power spectra of ITMX and ETMX -- payload part) (19117)

the frequency range in thir line is 0.07-0.11 Hz. Sorry.

CAL (XPcal)
dan.chen - 9:01 Wednesday 08 December 2021 (19122) Print this report
TCam check

Date: 2021/12/7

Members: Dan Chen, Yohei Nishino

We took some pictures of XTcam.
Before/after the work, we checked ETMX pictures to be sure the TCam alignment did not change because of the work.

In this time we used new LED lights (‎SG-01TB) which will be installed later in Tx/Rx module to illuminate ETM.

Picture 1: before
Picture 2: after

Other pictures, including TCam: https://dac.icrr.u-tokyo.ac.jp/KAGRA/DAWG/CAL/pic/Xend/211207

Images attached to this report
VIS (EX)
ryutaro.takahashi - 22:11 Tuesday 07 December 2021 (19121) Print this report
Finalization of EXV chamber

[Takahashi, Sato]

We went to the finalization of the chamber.

  1. Closed the feedthrough flange indicated by the red arrow in the picture.
  2. Cleaned the inside of the cross tube with the vacuum cleaner.
  3. The polarity of the F0 magnet had been recorded in the logbook. The center yoke is S.
Images attached to this report
VIS (EX)
ryutaro.takahashi - 21:20 Tuesday 07 December 2021 (19120) Print this report
Comment to ETMX BF V2 LVDT signal became noisy after the work on 30 of November (19107)

[Takahashi, Sato]

We investigated the noise in the BF V2 LVDT. In the work on 30 Nov., we moved the secondary coils. So the cable connection around the secondary coils might be changed. The resistances of the LVDT coils were the same as the measurement on 25 Nov. within 1 ohm, and the resistances were not changed when we shook the cables to the coils. The V2 spectrum itself is not so different from the V2 and V3 spectra. The increased noise may be due to the present DC level of V2 output  (12600 count) being much larger than the previous DC level (-1400 count).

Images attached to this comment
VIS (IX)
ryutaro.takahashi - 20:46 Tuesday 07 December 2021 (19119) Print this report
Comment to ITMX GAS LVDT signals seemed strange. (18789)

[Takahashi, Sato]

To investigate the noisy LVDT signals, we replaced the LVDT card for GAS filters (F1, F2, F3, BF). The modified card showed the noise floor higher than the original card's one. The phase and gain of the modified card have not been adjusted. In another case, the main (interface) card had a noisy output. We will try to replace the main card as well.

Images attached to this comment
VIS (IX)
masahide.tamaki - 20:21 Tuesday 07 December 2021 (19117) Print this report
Health check (checking the power spectra of ITMX and ETMX -- payload part)

Related  work: klog19107

I checked the spectra of ITMX and ETMX payload (as for tower part, please see Ushiba-san's work).

In MN DAMP Y spectrum, there were three strange peaks around 0.7-1.1 Hz (surrouded by blue circle in Fig1).
*Solid line and Dashed line show the spectrum after/before finalization work.

With Ushibasan's advice, I checked the spectrum of BF and found that there were peaks at the same frequency (Fig2). So this may affects to peaks shown in Fig1.
Seeing the sensor spectrum, there are similarly peaks for H1, H2, H3 (Fig3).  So we can say L and T are well balanced.

Besides this (ETMX IM and MN, ITMX IM), there was no problem.

Images attached to this report
Comments to this report:
masahide.tamaki - 9:03 Wednesday 08 December 2021 (19123) Print this report

the frequency range in thir line is 0.07-0.11 Hz. Sorry.

DetChar (General)
hirotaka.yuzurihara - 20:07 Tuesday 07 December 2021 (19118) Print this report
DQR : testing the command line of igwn-alert

Previous report : DQR : got the account on SciMMA auth system

Probably the trouble on the authentification was solved. I'm testing the command line of igwn-alert.

To receive the alert, we need to subscribe the topic (similar to channel) for each pipeline. Although I added several topics, it didn't work.... I will contact the developper. The detail is summarized on wiki page.

VIS (IX)
takaaki.yokozawa - 19:15 Tuesday 07 December 2021 (19116) Print this report
Acceptance check : IX oplevs
I tried to evaluate the spectrum with and without oplev beams.

Before of this work, I changed the D-Sub 15 pin for the length sensing QPD from hard one to soft one(Misumi order made)
And I closed the hole for the cables by aluminium. (See Fig.3 in picture)

After that I measured the oplev spectrum with turning off the oplev beams. (See Fig.4.)

Then, I tried to measure the spectrum with oplev beam, the suspension state was safe but this measurement was soon after the offloading work by Yasui-san and Ushiba-san, the beams were injected to QPDs (But not center), this spectrum is very temporal one.

Finally, I centered the length sensing QPDs (MNL, MNT, PFL).
Images attached to this report
VIS (IMM)
takaaki.yokozawa - 17:28 Tuesday 07 December 2021 (19115) Print this report
Check the optics for the IMMT1 oplev
[Kenta, Yokozawa]

We checked the optics for the IMMT1 oplev and we confirmed that oplev beam detected the QPD for IMMT1 tilt oplev.
Images attached to this report
Comments to this report:
takaaki.yokozawa - 12:50 Wednesday 08 December 2021 (19128) Print this report
One day trend (Sum value was about 1,100 from Yesterday, but we didn't perform centering yesterday...)
Images attached to this comment
VIS (IMM)
takaaki.yokozawa - 17:23 Tuesday 07 December 2021 (19114) Print this report
Recovery of the IMMT2 oplev
[Kenta, Yokozawa]

We performed the recovery of the IMMT2 oplev for the work of next week.

1. Check the tilt leveler around the place of setting pylon, and cleaning
2. Temporally set the pylon by checking the past klog15254 (Fig.1.)
3. We tried to search the good beam path without moving the optics on the IMMT oplev table, but we couldn't find the good alignment.
4. So we reset the position of optics to be better settings (Fig.3.)
- Set injection optics close to the viewport and set the QPD far from the viewport
- Because we don't need to consider data taking of the IMMT2 trans signal, so we changed the position, but if there is serious problem, please let us know
5. Check the beam path on the IMMT2 mirror and reflection mirror to be centering (Fig.2.)
6. Now, we can use IMMT2 oplev, and we centered it around 16:00 today.
7. Some dumper and gige camera was placed near the table (Fig.4.)

Images attached to this report
VIS (BS)
fabian.arellano - 15:46 Tuesday 07 December 2021 (19104) Print this report
Comment to In-vacuum cable health check at outer side of flanges. (18191)

With Terrence.

Summary: we corroborated that the new adapter circuit for the F1 LVDT (entry 18964) changes the polarity of the readout. We should change the polarity of the primary coil within the adapter box in order to fix this.

Terrence reported recently that the polarity of the actuation-LVDT-readout system was apposite to what we had before, so I checked.

First I corroborated that negative actuation produced positive readout from the LVDT. This is not what we want. Then I removed the adapter circuit  (entry 18964) and corroborated that the correct polarity was recovered.

We would like to set the polarity in hardware, not in software. The appropriate way of doing it would be to exchange the places of pins 2 and 7 inside of the adapter circuit box. Pins 2 and 7 connect to the LVDT primary coil, which is healthy and is used as a readout with the adapter circuit.

FCL (Air)
shinji.miyoki - 15:24 Tuesday 07 December 2021 (19113) Print this report
Comment to 20.5 -> 19.5 for the temp setting of precision cooler near SR3 (19105)

Because there seems to be no serious rapid change of temp, I changed from 19.50 -> 19.00. (18.00 is the minimum setting temp for this precision cooler)

By the way, the hot water was abandoned from this precision cooler to the gutter in the corner station as the result of the heat exchange from the air to water. However, I noticed that this hot water heat up again the air in the corner station even if the hot water is mixed with cold water in the gutter, and even if most gutters were covered with concrete lids. I will ask Hayakawa-kun to abandon the hot water not in the same area but in the pre-room at least.

CRY (Cryostat EX)
takayuki.tomaru - 15:13 Tuesday 07 December 2021 (19112) Print this report
Temperature Monitor Trouble of 4K-Ref.2
Since the temperature of the cold-head at 4K-Ref.2 were not displayed on the CryoCon temperature monitor, I checked this problem. Finally I concluded that this trouble came from noise on the cryostat. After I disconnected and reconnected the burndy connector on the 4K-Ref.2, the temperature display was recovered. Cabling for the thermometers and CryoCon were no problem.
PEM (Center)
takaaki.yokozawa - 15:00 Tuesday 07 December 2021 (19111) Print this report
Acceptance check: IX TCam
[DanChen Yokozawa]

As one of the acceptance check, we checked the TCam.
First, we illuminated from the TM oplev viewport by the LED light.
Then we tuned the rotation and took picture.
We confirmed that we can take data without IR/GR laser and obsolete illuminator
Images attached to this report
VAC (IXA)
shinji.miyoki - 14:55 Tuesday 07 December 2021 (19110) Print this report
No lost item in IXA tank

[miyoki, chendan]

We inspect the inside of IXA tank (actually the area above the optical table) to make assure there is no nasty thing inside IXA. According to our eyeball inspection. no nasty thing inside. 

By the way, maybe as you know, there is a small window (ICF70) on the IXA tank (-Y, +X) as in the attached photo. 

Images attached to this report
Search Help
×

Warning

×